OpenStudy (anonymous):

Which are real zeroes of this function? x3 + 2x2 – 9x – 18

OpenStudy (amoodarya):

\[x^3 + 2x^2 – 9x – 18=(x^3 + 2x^2 ) -(9x+18 ) = \] \[(x^3 + 2x^2 ) -(9x+18 ) = x^2(x+2) -9(x+2)\] \[x^2(x+2) -9(x+2) = (x+2)(x^2-9)\] can you go on ?